OpenStudy (tayaiv):

Triangle ABC has side lengths: √6, √2, and 2√2 units. If the base of the triangle has a length of √6, what is the measure of the base angle? I'm not entirely sure which base angle the question is talking about. Thanks!

OpenStudy (dayakar):



OpenStudy (dayakar):

\[a= \sqrt{6}, b= \sqrt{2}, c= 2\sqrt{2}\]

OpenStudy (lochana):

this is a right triangle. because side lengths obey pythsgoras theorem (2root(2))^2 = 6 + 2 8 = 8

OpenStudy (lochana):



OpenStudy (dayakar):

\[\cos A = \frac{ b ^{2}+c ^{2}-a ^{2} }{ 2bc}\]

OpenStudy (dayakar):

u r right I used above formula

OpenStudy (lochana):

sorry . it is cosx. not sinx. so cosx = root(3)/2 x = 30 degrees :)

OpenStudy (tayaiv):

Thank you so much!! @lochana @dayakar